landon:

On Monday, a company stock closed at a price of $30.80 per share. On Tuesday, it increased by $1.20 per share. On Wednesday, it fell by 6.25%. On Thursday, it fell by another 4%, and on Friday, it increased by 2.5%. What is the net change in the stock price from Monday to Friday? A. -$1.28 B. -$0.40 C. $0.40 D. $2.00

dude:

Okay so you want to find out the change per day Monday = 30.80 Tuesday = 1.20 1. For Monday-Tuesday you want to add up the values (Because it was a simple $ change) Monday + Tuesday = Net for Tuesday 2. For Wednesday, you want to calculate what 6.25% is of the total for Monday and Tuesday Net for Wednesday= Net for Tuesday \(\times\) 0.625 [0.625 is the decimal of the percent, you cannot multiply percents] 3. From then, calculate 4% of Wednesday's price and subtract it Net for Thursday = Net for Wednesday \(\times\) 0.04 [Same as previous step, 0.04 is 4% as a decimal] 4. For Friday, calculate 2.5% of Thursdays price and add it Answer = Net for Thursday \(\times\) 0.25 [2.5% as a decimal]